OpenStudy (anonymous):

6ySqrT(7)*2SqrT(y)=12ySqrT(7y) Am I right?

OpenStudy (espex):

Yes.\[6y\sqrt{7}*2\sqrt{y}=12y\sqrt{7y}\]

OpenStudy (anonymous):

Thank You!(:

OpenStudy (espex):

You're welcome.